# DRL-Based Federated Self-Supervised Learning for Task Offloading and Resource Allocation in ISAC-Enabled Vehicle Edge Computing

### v1

@@ Introduction
$\eta_{q_{0}}=\frac{1}{NS^{max}}\sum_{n=1}^{N}\sum_{t=1}^{S^{max}}\left(\frac{N_{t}^{n,off}}{N_{t}^{n,total}}\times 100\%\right).$ (33)
@@
The problem of task offloading and resource allocation can be addressed with a DRL algorithm. To this end, we will utilize DRL to find the optimal task offloading and resource allocation scheme. The SAC algorithm is chosen due to its robustness and efficiency in handling continuous action spaces and its ability to balance exploration and exploitation effectively. Compared to other DRL algorithms, SAC introduces an entropy term into the objective function to encourage BS to continuously explore new policies. The presence of this entropy term effectively prevents BS from prematurely converging to a local optimal solution during policy iteration, thereby preventing training failure.
@@
For the decision-making process in each slot $t$, it includes state, action, reward, and solution. We define the state, action, reward, and solution in slot $t$ as follows:
@@
$s_{t}=\left(\mathbb{G},\mathbb{V},\mathbb{N}\right).$ (34)
@@
$a_{t}=\left(\mathbb{P},\mathbb{F},\mathbb{Q}\right).$ (35)
@@
Our objective is to minimize energy consumption, overload amount, and the remaining local iterations in the buffer within each slot. Thus, in slot $t$, the reward consists of three negative components: the total energy consumption $E_{t}^{n,total}$, the overload amount $O_{t}^{n}$, and the remaining local iterations $B_{t}^{n}$ at the end of slot $t$. The reward function can be calculated as:
$r_{t}=-\sum_{n=1}^{N}{\left(\vartheta_{1}E_{t}^{n,total}+\vartheta_{2}O_{t}^{n}+\vartheta_{3}B_{t}^{n}\right)},$ (36)
@@
The SAC algorithm bases on the maximum entropy principle, and it involves five networks: one actor network $\varsigma$, two critic networks $\xi_{1}$ and $\xi_{2}$, and two target networks $\varpi_{1}$ and $\varpi_{2}$. Following, we will introduce the complete SAC-based Federated SSL solution for task offloading and resource allocation.
@@
Step 2, task offloading and training: In episode $e$, before slot $t$ starts, determine the relationship between round $r$ and slot $t$ according to Eq. (1) firstly. BS stores the parameters of a global model $\theta_{{r-1}}^{\text{global}}$ aggregated in previous round $r-1$, along with the five networks of SAC algorithm. If in episode $e$, slot $1$, the parameter of five networks come from episode $e-1$, slot $S^{max}$. Vehicles also store $\mathcal{Z}$ image data, denoted as local data, and a local model which has the same structure as global model. Then, BS employs SAC algorithm to allocate $p_{t,R}^{n}$, $p_{t,B}^{n}$, $f_{t,L}^{n,comp}$, and $q_{t}^{n}$ for vehicle $n$. Next, vehicle $n$ downloads the global model $\theta_{{r-1}}^{\text{global}}$ from BS and assigns the parameters of global model to local model $\theta_{{r}}^{\text{local},n}$, along with the transmission power $p_{t,R}^{n}$, $p_{t,B}^{n}$, $f_{t,L}^{n,comp}$, and $q_{t}^{n}$ allocated by BS for vehicle $n$. Based on the downloaded CPU frequency $f_{t,L}^{n,comp}$, the total training time $T$, and true training time $T^{\prime}$, $N_{t}^{n}$ and ${N_{t}^{n}}^{\ast}$ can be calculated by Eq. (17) and Eq. (18). Then, conducting task allocating to obtain $N_{t,L}^{n}$ and $N_{t,R}^{n,off}$ according to Section 4.1. If the downloaded assignment ratio $q_{t}^{n}$ is greater than the offloading threshold $q_{0}$, the vehicle $n$ immediately sends an offloading request to RSU with the downloaded transmission power $p_{t,R}^{n}$. This request includes local data $\left\{x_{r}^{1,n},\ x_{r}^{2,n},\ \ldots x_{r}^{i,n},\ldots x_{r}^{\mathcal{Z},n}\right\}$, local model $\theta_{{r}}^{\text{local},n}$, and assignment ratio $q_{t}^{n}$.
@@
Meanwhile, vehicle $n$ also performs local training with local data and local model for $N_{t,L}^{n}$ iterations. The training process is Federated SSL. For each iteration, first, input each local data $x_{r}^{i,n}$, $x_{r}^{i,n}\in\left\{x_{r}^{1,n},\ x_{r}^{2,n},\ \ldots x_{r}^{i,n},\ldots x_{r}^{\mathcal{Z},n}\right\}$, into two different data augmentation methods, $\pi_{1}(\cdot)$ and $\pi_{2}(\cdot)$. The remaining images $x_{r}^{j,n},j\in\left[1,\mathcal{Z}\right]\ \text{and}\ j\neq i$ are denoted as negative samples. Feeding augmented images $x_{r}^{i,n}$ and $x_{r}^{j,n}$ into a function $\chi$, which is composed of an encoder, two dense neutral networks and a ReLU function. The output of the $\chi$ includes anchor sample $q_{r}^{i,{n}}$, positive sample $k_{{r}}^{i,{n}}$, and encoded negative samples $k_{{r}}^{j,{n}}$, respectively, i.e.,
$q_{r}^{i,n}=\chi\left\{\pi_{1}\left(x_{r}^{i,n}\right)\right\},\ i\in\left[1,\mathcal{Z}\right],$ (37)
$k_{r}^{i,n}=\chi\left\{\pi_{2}\left(x_{r}^{i,n}\right)\right\},\ i\in\left[1,\mathcal{Z}\right],$ (38)
$k_{r}^{j,n}=\chi\left\{x_{r}^{j,n}\right\},\ j\in\left[1,\mathcal{Z}\right],\ \text{and}\ j\neq i.$ (39)
@@
$\begin{split}\mathcal{L}_{{q_{{r}}^{i,n}}|\tau={\tau_{1}}}=\frac{\exp{\left(\frac{q_{{r}}^{i,n}\cdot k_{{r}}^{i,n}}{\tau}\right)}}{\exp{\left(\frac{q_{{r}}^{i,n}\cdot k_{{r}}^{i,n}}{\tau}\right)}+\sum_{j=1}^{\mathcal{K}}\exp{\left(\frac{q_{{r}}^{i,n}\cdot k_{{r}}^{j,n}}{\tau}\right)}},\end{split}$ (40)
where $\mathcal{K}$ is the number of negative samples, and $\tau$ is temperature hyper-parameters [61]. Thus, the dual temperature loss used by FLSimCo can be calculated as
$\begin{split}\mathcal{L}_{q_{{r}}^{i,n}}^{\text{DT}}=\text{sg}\left[\frac{1-\mathcal{L}_{{q_{{r}}^{i,n}}|\tau={\tau_{2}}}}{1-\mathcal{L}_{{q_{{r}}^{i,n}}|\tau={\tau_{1}}}}\right]\times\log\mathcal{L}_{{q_{{r}}^{i,n}}|\tau={\tau_{1}}},\end{split}$ (41)
where $\text{sg}[\cdot]$ indicates the stop gradient. We use the Stochastic Gradient Descent (SGD) method to update the local model [62], i.e.,
@@
$\theta_{{r,L}}^{\text{local},n}=\theta_{{r}}^{\text{local},n}-\eta^{r}\left[\nabla_{q_{{r}}^{i,n}}\left(\frac{1}{\mathcal{Z}}\sum_{i=1}^{\mathcal{Z}}\mathcal{L}_{q_{{r}}^{i,n}}^{\text{DT}}\right)\right],$ (42)
@@
Step 3, uploading: After receiving all trained model sent by RSU, vehicles upload trained models $\left\{\theta_{r,L}^{local,1},\theta_{r,L}^{local,2},\cdots\theta_{r,L}^{local,n},\cdots\theta_{r,L}^{local,N}\right\}$, $\left\{\theta_{r,R}^{local,1},\theta_{r,R}^{local,2},\cdots\theta_{r,R}^{local,n},\cdots\theta_{r,R}^{local,N}\right\}$, velocities $\left\{v_{t+1}^{1},v_{t+1}^{2},\cdots v_{t+1}^{n},\cdots v_{t+1}^{N}\right\}$ for slot $t+1$, total energy consumption $\left\{E_{t}^{1,total},E_{t}^{2,total},\cdots E_{t}^{n,total},\cdots E_{t}^{N,total}\right\}$, and unprocessed training iterations $\left\{B_{t}^{1},B_{t}^{2},\cdots B_{t}^{n},\cdots B_{t}^{N}\right\}$ to BS with the downloaded transmission power $\left\{p_{t,B}^{1},p_{t,B}^{2},\cdots p_{t,B}^{n},\cdots p_{t,B}^{N}\right\}$, respectively.
Step 4, aggregating and updating: Upon receiving the models from all vehicles, the BS aggregates received models followed as:
$\theta_{r}^{\text{global}}=\frac{\sum_{n=1}^{N}\left(\theta_{r,L}^{\text{local},n}+\theta_{r,R}^{\text{local},n}\right)}{2N},$ (43)
@@
$J(\pi_{\varsigma})=E_{(s_{t},a_{t})\sim\pi_{\varsigma}}\\ \left[\gamma^{t-1}r_{t}\right.\left.+\beta H(\pi_{\varsigma}\left(\cdot|s_{t}\right))\right],$ (44)
@@
$\beta^{\ast}=\arg\max_{\beta}J(\pi_{\varsigma}^{\ast}).$ (45)
@@
$\begin{split}\nabla_{\beta}J_{\beta}=-\nabla_{\beta}\ E_{a_{m}^{\prime}\sim\pi_{\varsigma}}\left[\beta\log\pi_{\varsigma}(a_{m}^{\prime}|s_{m})+\beta\widetilde{H}\right],\end{split}$ (46)
@@
Next, as shown by the yellow line in Fig. 3, we start to update the actor network $\varsigma$. Firstly, feeding $(s_{m},a_{m}^{\prime})$ into the two critic networks $\xi_{1}$ and $\xi_{2}$, respectively, yielding the corresponding action value functions $Q_{\xi_{1}}(s_{m},a_{m}^{\prime})$ and $Q_{\xi_{2}}(s_{m},a_{m}^{\prime})$. The smaller of these two values is defined as $Q_{\xi}(s_{m},a_{m}^{\prime})$. Thus, by the SGD algorithm, the actor network $\varsigma$ is updated as follows:
$\begin{split}\nabla_{\varsigma}\mathcal{J}_{\varsigma}&=\nabla_{\varsigma}\left[-\frac{1}{M}\sum_{m=1}^{M}\left(\beta\log\pi_{\varsigma}\left(a_{m}^{\prime}\middle|s_{m}\right)\right)^{2}\right]\\ &+\nabla_{\varsigma}\left[-\frac{1}{M}\sum_{m=1}^{M}\left(f_{\varsigma}(o;s_{m})\times\nabla_{a_{m}^{\prime}}\left(Q_{\xi}(s_{m},a_{m}^{\prime})\right.\right.\right.\\ &\left.\left.\left.-\beta\log\pi_{\varsigma}\left(a_{m}^{\prime}\middle|s_{m}\right)\right)\right)^{2}\right],\end{split}$ (47)
@@
$\begin{split}\nabla_{\xi_{1}}&J_{\xi_{1}}=\nabla_{\xi_{1}}Q_{\xi_{1}}(s_{m},a_{m})\times\left[Q_{\xi_{1}}(s_{m},a_{m})-r_{m}\right.\\ &\left.+\gamma[-\beta\log\pi_{\varsigma}^{m+1}(a_{m+1}|s_{m+1})+Q_{\varpi}(s_{m+1},a_{m+1})]\right],\end{split}$ (48)
$\begin{split}\nabla_{\xi_{2}}&J_{\xi_{2}}=\nabla_{\xi_{2}}Q_{\xi_{2}}(s_{m},a_{m})\times\left[Q_{\xi_{2}}(s_{m},a_{m})-r_{m}\right.\\ &\left.+\gamma[-\beta\log\pi_{\varsigma}^{m+1}(a_{m+1}|s_{m+1})+Q_{\varpi}(s_{m+1},a_{m+1})]\right].\end{split}$ (49)
@@
$\varpi_{1}:=\omega\xi_{1}+(1-\omega)\varpi_{1},$ (50)
$\varpi_{2}:=\omega\xi_{2}+(1-\omega)\varpi_{2},$ (51)
@@
This experiment is entirely based on the scenario described in Section 3, and utilizes Python 3.10 for simulation. The actor network and both critic networks are implemented as four-layer fully connected DNNs, each featuring two hidden layer sand containing 512 neurons. The following training simulation results represent the average of three experiments, and the environmental parameters and SAC algorithm hyper-parameters used in this paper are shown in Table 2.
@@
(a) Energy consumption
(b) Calculation comparison
(c) Reward comparison
Fig. 4: Comparison between different DRL algorithms
@@
Fig. 4 shows a comparison of different methods with five vehicles taking part in the training. The training methods include four DRL methods: SAC, DDPG, TD3, and PPO, as well as a method always transmitting with maximum power. We assess these methods in terms of energy consumption, computation times, and reward. It is obvious that the method of transmitting with maximum power results in the highest energy consumption, highest computation times and lowest reward. In contrast, other DRL methods achieve varying degrees of reducing energy consumption. Next, we will focus on comparing the four DRL methods.
@@
Regarding energy consumption, as shown in Fig. 4(a), PPO algorithm fails to effectively reduce energy consumption, with its energy consumption fluctuating around $(43.69\pm 1.24)J$. In contrast, the energy consumption of the SAC, DDPG, and TD3 algorithms shows a rapid decline and gradually converges to a relatively stable value. Specifically, the SAC algorithm exhibits the fastest convergence speed and is more stable compared to other two algorithms, with the final energy consumption stabilizing at $(2.18\pm 1.02)J$. The inset provides detailed information for the episodes range $\left[1,500\right]$. It is evident from the inset that TD3 algorithm exhibits less fluctuation than DDPG algorithms but DDPG algorithm can explore lower energy consumption values. Ultimately, the energy consumption of the DDPG algorithm stabilizes at $(8.95\pm 6.87)J$, while that of the TD3 algorithm stabilizes at $(6.24\pm 3.54)J$.
@@
Overall, the SAC algorithm demonstrates the lowest energy consumption. Despite its final reward being slightly lower than that of the DDPG and TD3 algorithms, SAC algorithm converges rapidly and stably across all three metrics. This effectiveness is primarily due to the capability of SAC algorithm to balance exploration and exploitation during policy updates, enhancing convergence efficiency. Consequently, SAC algorithm proves to be the most effective in achieving our goal of minimizing energy consumption, making it the best-performing algorithm.
In Fig. 7, we compare the loss function, Top1 accuracy, and Top5 accuracy of the global model of FLSimCo (without offloading) and our method (with offloading) under different numbers of vehicles. It is obvious that the loss function gradually decreases and converges, while the accuracy steadily increases and converges. Specifically, as illustrated in Fig. 7(a), when comparing the cases of 5 vehicles (with 5 models aggregated per round) and 10 vehicles (with 10 models aggregated per round) without task offloading, we observe that with 10 vehicles, the loss function fluctuates more and converges more slowly. This indicates that as the number of vehicles increases, the model diversity also increases, negatively impacting the effectiveness of federated aggregation. However, by offloading tasks to the RSU for training, we obtain models trained by the RSU, which are then sent back to the BS for aggregation along with the locally trained models. Therefore, with 5 vehicles and task offloading, the final number of models is doubled, resulting in 10 local models aggregated per round, which is the same as the number of models aggregated by the BS with 10 vehicles without task offloading. From the figure, we can see that the 5 vehicles with task offloading (10 models in total) not only outperform the 10 vehicles without task offloading but also the 5 vehicles without task offloading. This is because offloading some tasks to the RSU allows the completion of all scheduled iterations, and more local iterations help mitigate the negative impact of model diversity on federated aggregation. Fig. 7(b) and Fig. 7(c) show that in the first 60 episodes, the fewer models aggregated per round, the faster the accuracy improves. However, as the number of episodes increases, this diversity gradually diminishes, resulting in the accuracy of global with 5 and 10 vehicles without task offloading gradually converging. Additionally, when 5 vehicles participate in Federated SSL and offload partial iterations task to RSU, the number of models used in the final aggregation is 10. During the first 60 episodes, the accuracy is not affected by model diversity. Instead, it increases rapidly and steadily. In the later stages of training, due to more iterations per round, the accuracy improvement is significantly better than others. This indicates that a greater number of iterations mitigates the negative effects of model diversity during aggregation.
@@
In Fig. 8, we compare the energy consumption and reward obtained by the SAC algorithm with different numbers of vehicles. From Fig. 8(a) and Fig. 8(b), it can be observed that the energy consumption curve shows a decreasing trend and gradually stabilize and the reward function shows a increasing trend and gradually stabilize. Furthermore, as the number of vehicles increases, the number of episodes required to achieve a stable state also increases. This indicates that in multi-vehicle scenarios, task offloading and resource allocation become more complex, leading to a greater number of episodes being required to achieve convergence to a stable state. Overall, as the number of vehicles participating in the training gradually increases, the total energy consumption of the system also increases, while the reward values decrease. Furthermore, through these two figures, we can also clearly see the superiority of the SAC algorithm in terms of convergence speed and stability.

### v2

@@ Introduction
$\eta_{q_{0}}=\frac{1}{NS^{max}}\sum_{n=1}^{N}\sum_{t=1}^{S^{max}}\left(\frac{N_{t}^{n,off}}{N_{t}^{n,total}}\times 100\%\right)$ (33)
@@
The problem of task offloading and resource allocation can be addressed with a DRL algorithm. To this end, we will utilize DRL to find the optimal task offloading and resource allocation scheme. The SAC algorithm is chosen due to its robustness and efficiency in handling continuous action spaces and its ability to balance exploration and exploitation effectively. SAC is an off-policy DRL algorithm designed to enhance the stability of learning and the exploration of the policy. It encourages exploration through the maximum entropy framework, preventing premature convergence to suboptimal solutions. Additionally, SAC uses DQN to reduce overestimation bias in value function approximation, further improving the stability. Compared to other DRL algorithms, SAC introduces an entropy term into the objective function to encourage BS to continuously explore new policies. The presence of this entropy term effectively prevents BS from prematurely converging to a local optimal solution during policy iteration, thereby preventing training failure.
@@
For the decision-making process in each slot $t$, it includes state, action, reward, and solution. We define the state, action, reward, and solution in slot $t$ as follows
@@
$s_{t}=\left(\mathbb{G},\mathbb{V},\mathbb{N}\right)$ (34)
@@
$a_{t}=\left(\mathbb{P},\mathbb{F},\mathbb{Q}\right)$ (35)
@@
Our objective is to minimize energy consumption, overload amount, and the remaining local iterations in the buffer within each slot. Thus, in slot $t$, the reward consists of three negative components: the total energy consumption $E_{t}^{n,total}$, the overload amount $O_{t}^{n}$, and the remaining local iterations $B_{t}^{n}$ at the end of slot $t$. The reward function can be calculated as
$r_{t}=-\sum_{n=1}^{N}{\left(\vartheta_{1}E_{t}^{n,total}+\vartheta_{2}O_{t}^{n}+\vartheta_{3}B_{t}^{n}\right)}$ (36)
@@
The SAC algorithm bases on the maximum entropy principle, and it involves five networks: one actor network $\varsigma$, two critic networks $\xi_{1}$ and $\xi_{2}$, and two target networks $\varpi_{1}$ and $\varpi_{2}$. At the start of the scheme, the BS holds the parameters of the global model, while each vehicle retains a local model that has the same structure as the global model. Following, we will introduce the complete SAC-based federated SSL solution for task offloading and resource allocation.
@@
Step 2, task offloading and training: In episode $e$, before slot $t$ starts, determine the relationship between round $r$ and slot $t$ according to Eq. (1) firstly. BS stores the parameters of a global model $\theta_{{r-1}}^{\text{global}}$ aggregated in previous round $r-1$, along with the five networks of SAC algorithm. If in episode $e$, slot $1$, the parameter of five networks come from episode $e-1$, slot $S^{max}$. Vehicles also store $\mathcal{Z}$ image data, denoted as local data, and a local model which has the same structure as global model. Then, BS employs SAC algorithm to allocate $p_{t,R}^{n}$, $f_{t,L}^{n,comp}$, and $q_{t}^{n}$ for vehicle $n$. Next, vehicle $n$ downloads the global model $\theta_{{r-1}}^{\text{global}}$ from BS and assigns the parameters of global model to local model $\theta_{{r}}^{\text{local},n}$, along with the transmission power $p_{t,R}^{n}$, $f_{t,L}^{n,comp}$, and $q_{t}^{n}$ allocated by BS for vehicle $n$. Based on the downloaded CPU frequency $f_{t,L}^{n,comp}$, the total training time $T$, and true training time $T^{\prime}$, $N_{t}^{n}$ and ${N_{t}^{n}}^{\ast}$ can be calculated by Eq. (17) and Eq. (18). Then, conducting task allocating to obtain $N_{t,L}^{n}$ and $N_{t,R}^{n,off}$ according to Section 4.1. If the downloaded assignment ratio $q_{t}^{n}$ is greater than the offloading threshold $q_{0}$, the vehicle $n$ immediately sends an offloading request to RSU with the downloaded transmission power $p_{t,R}^{n}$. This request includes local data $\left\{x_{r}^{1,n},\ x_{r}^{2,n},\ \ldots x_{r}^{i,n},\ldots x_{r}^{\mathcal{Z},n}\right\}$, local model $\theta_{{r}}^{\text{local},n}$, and assignment ratio $q_{t}^{n}$.
Meanwhile, vehicle $n$ also performs local training with local data and local model for $N_{t,L}^{n}$ iterations. The training process is a federated SSL. For each iteration, first, input each local data $x_{r}^{i,n}$, $x_{r}^{i,n}\in\left\{x_{r}^{1,n},\ x_{r}^{2,n},\ \ldots x_{r}^{i,n},\ldots x_{r}^{\mathcal{Z},n}\right\}$, into two different data augmentation methods, $\pi_{1}(\cdot)$ and $\pi_{2}(\cdot)$. The function $\pi_{1}(\cdot)$ applies a horizontal flip to the image with a 50% likelihood, and converts the image to grayscale with a 20% likelihood. Meanwhile, $\pi_{2}(\cdot)$ adjusts the brightness of image, saturation, and hue with an 80% likelihood, and then converts the image to grayscale with a 40% likelihood. These two different augmentations are adopted in the FLSimCo, so the features of image can be efficiently utilized.
@@
The remaining images $x_{r}^{j,n},j\in\left[1,\mathcal{Z}\right]\ \text{and}\ j\neq i$ are denoted as negative samples. Feeding augmented images $x_{r}^{i,n}$ and $x_{r}^{j,n}$ into a function $\chi$, which is composed of an encoder, two dense neutral networks and a ReLU function. The output of the $\chi$ includes anchor sample $q_{r}^{i,{n}}$, positive sample $k_{{r}}^{i,{n}}$, and encoded negative samples $k_{{r}}^{j,{n}}$, respectively, i.e.,
$q_{r}^{i,n}=\chi\left\{\pi_{1}\left(x_{r}^{i,n}\right)\right\},\ i\in\left[1,\mathcal{Z}\right]$ (37)
$k_{r}^{i,n}=\chi\left\{\pi_{2}\left(x_{r}^{i,n}\right)\right\},\ i\in\left[1,\mathcal{Z}\right]$ (38)
$k_{r}^{j,n}=\chi\left\{x_{r}^{j,n}\right\},\ j\in\left[1,\mathcal{Z}\right],\ \text{and}\ j\neq i$ (39)
@@
$\begin{split}\mathcal{L}_{{q_{{r}}^{i,n}}|\tau={\tau_{1}}}=\frac{\exp{\left(\frac{q_{{r}}^{i,n}\cdot k_{{r}}^{i,n}}{\tau}\right)}}{\exp{\left(\frac{q_{{r}}^{i,n}\cdot k_{{r}}^{i,n}}{\tau}\right)}+\sum_{j=1}^{\mathcal{K}}\exp{\left(\frac{q_{{r}}^{i,n}\cdot k_{{r}}^{j,n}}{\tau}\right)}}\end{split}$ (40)
where $\mathcal{K}$ is the number of negative samples, and $\tau$ is temperature hyper-parameter [54]. Thus, the dual temperature loss used by FLSimCo can be calculated as
$\begin{split}\mathcal{L}_{q_{{r}}^{i,n}}^{\text{DT}}=\text{sg}\left[\frac{1-\mathcal{L}_{{q_{{r}}^{i,n}}|\tau={\tau_{2}}}}{1-\mathcal{L}_{{q_{{r}}^{i,n}}|\tau={\tau_{1}}}}\right]\times\log\mathcal{L}_{{q_{{r}}^{i,n}}|\tau={\tau_{1}}}\end{split}$ (41)
where $\text{sg}[\cdot]$ indicates the stop gradient. $\tau_{1}$ and $\tau_{2}$ are different values of temperature hyper-parameter $\tau$. We use the Stochastic Gradient Descent (SGD) method to update the local model [55], i.e.,
@@
$\theta_{{r,L}}^{\text{local},n}=\theta_{{r}}^{\text{local},n}-\eta^{r}\left[\nabla_{q_{{r}}^{i,n}}\left(\frac{1}{\mathcal{Z}}\sum_{i=1}^{\mathcal{Z}}\mathcal{L}_{q_{{r}}^{i,n}}^{\text{DT}}\right)\right]$ (42)
@@
Step 3, uploading: After receiving all trained model sent by RSU, vehicles upload trained models $\left\{\theta_{r,L}^{local,1},\theta_{r,L}^{local,2},\cdots\theta_{r,L}^{local,n},\cdots\theta_{r,L}^{local,N}\right\}$, $\left\{\theta_{r,R}^{local,1},\theta_{r,R}^{local,2},\cdots\theta_{r,R}^{local,n},\cdots\theta_{r,R}^{local,N}\right\}$, velocities $\left\{v_{t+1}^{1},v_{t+1}^{2},\cdots v_{t+1}^{n},\cdots v_{t+1}^{N}\right\}$ for slot $t+1$, total energy consumption $\left\{E_{t}^{1,total},E_{t}^{2,total},\cdots E_{t}^{n,total},\cdots E_{t}^{N,total}\right\}$, and unprocessed training iterations $\left\{B_{t}^{1},B_{t}^{2},\cdots B_{t}^{n},\cdots B_{t}^{N}\right\}$ to BS, respectively.
Step 4, aggregating and updating: Upon receiving the models from all vehicles, the BS aggregates received models followed as
$\theta_{r}^{\text{global}}=\frac{\sum_{n=1}^{N}\left(\theta_{r,L}^{\text{local},n}+\theta_{r,R}^{\text{local},n}\right)}{2N}$ (43)
@@
$J(\pi_{\varsigma})=E_{(s_{t},a_{t})\sim\pi_{\varsigma}}\\ \left[\gamma^{t-1}r_{t}\right.\left.+\beta H(\pi_{\varsigma}\left(\cdot|s_{t}\right))\right]$ (44)
@@
$\beta^{\ast}=\arg\max_{\beta}J(\pi_{\varsigma}^{\ast})$ (45)
@@
$\begin{split}\nabla_{\beta}J_{\beta}=-\nabla_{\beta}\ E_{a_{m}^{\prime}\sim\pi_{\varsigma}}\left[\beta\log\pi_{\varsigma}(a_{m}^{\prime}|s_{m})+\beta\widetilde{H}\right]\end{split}$ (46)
@@
Next, as shown by the yellow line in Fig. 3, we start to update the actor network $\varsigma$. Firstly, feeding $(s_{m},a_{m}^{\prime})$ into the two critic networks $\xi_{1}$ and $\xi_{2}$, respectively, yielding the corresponding action value functions $Q_{\xi_{1}}(s_{m},a_{m}^{\prime})$ and $Q_{\xi_{2}}(s_{m},a_{m}^{\prime})$. The smaller of these two values is defined as $Q_{\xi}(s_{m},a_{m}^{\prime})$. Thus, by the SGD algorithm, the actor network $\varsigma$ is updated as follows
$\begin{split}\nabla_{\varsigma}\mathcal{J}_{\varsigma}&=\nabla_{\varsigma}\left[-\frac{1}{M}\sum_{m=1}^{M}\left(\beta\log\pi_{\varsigma}\left(a_{m}^{\prime}\middle|s_{m}\right)\right)^{2}\right]\\ &+\nabla_{\varsigma}\left[-\frac{1}{M}\sum_{m=1}^{M}\left(f_{\varsigma}(o;s_{m})\times\nabla_{a_{m}^{\prime}}\left(Q_{\xi}(s_{m},a_{m}^{\prime})\right.\right.\right.\\ &\left.\left.\left.-\beta\log\pi_{\varsigma}\left(a_{m}^{\prime}\middle|s_{m}\right)\right)\right)^{2}\right]\end{split}$ (47)
@@
$\begin{split}\nabla_{\xi_{1}}&J_{\xi_{1}}=\nabla_{\xi_{1}}Q_{\xi_{1}}(s_{m},a_{m})\times\left[Q_{\xi_{1}}(s_{m},a_{m})-r_{m}\right.\\ &\left.+\gamma[-\beta\log\pi_{\varsigma}^{m+1}(a_{m+1}|s_{m+1})+Q_{\varpi}(s_{m+1},a_{m+1})]\right]\end{split}$ (48)
$\begin{split}\nabla_{\xi_{2}}&J_{\xi_{2}}=\nabla_{\xi_{2}}Q_{\xi_{2}}(s_{m},a_{m})\times\left[Q_{\xi_{2}}(s_{m},a_{m})-r_{m}\right.\\ &\left.+\gamma[-\beta\log\pi_{\varsigma}^{m+1}(a_{m+1}|s_{m+1})+Q_{\varpi}(s_{m+1},a_{m+1})]\right]\end{split}$ (49)
@@
$\varpi_{1}:=\omega\xi_{1}+(1-\omega)\varpi_{1}$ (50)
$\varpi_{2}:=\omega\xi_{2}+(1-\omega)\varpi_{2}$ (51)
@@
This experiment is entirely based on the scenario described in Section 3, and utilizes Python 3.10 for simulation. The actor network and both critic networks are implemented as four-layer fully connected DNNs, each featuring two hidden layer sand containing 512 neurons. We set up a scenario with four intersections. In urban environments, the variety of object categories is relatively limited, and images of objects within the same category are frequently captured. Therefore, we chose the CIFAR-10 dataset, which contains 10 common categories with 5,000 distinct images per category.
The following training simulation results represent the average of three experiments, and the environmental parameters and SAC algorithm hyper-parameters used in this paper are shown in Table 2.
@@
(a) Energy consumption
(b) Calculation comparison
(c) Reward comparison
Fig. 4: Comparison between different DRL algorithms
@@
In Fig. 4, we show a comparison of different DRL methods with five vehicles taking part in each round of FL. The training methods include SAC, DDPG, TD3, and PPO. We assess these methods in terms of energy consumption, computation times, and reward. Regarding energy consumption, as shown in Fig. 4(a), PPO algorithm fails to effectively reduce energy consumption, with its energy consumption fluctuating around $(43.69\pm 1.24)J$. In contrast, the energy consumption of the SAC, DDPG, and TD3 algorithms shows a rapid decline and gradually converges to a relatively stable value. Specifically, the SAC algorithm exhibits the fastest convergence speed and is more stable compared to other two algorithms, with the final energy consumption stabilizing at $(2.18\pm 1.02)J$. The inset provides detailed information for the episodes range $\left[1,500\right]$. It is evident from the inset that TD3 algorithm exhibits less fluctuation than DDPG algorithms but DDPG algorithm can explore lower energy consumption values. Ultimately, the energy consumption of the DDPG algorithm stabilizes at $(8.95\pm 6.87)J$, while that of the TD3 algorithm stabilizes at $(6.24\pm 3.54)J$.
@@
Overall, the SAC algorithm demonstrates the lowest energy consumption. Although its final reward is slightly lower than those of the DDPG and TD3 algorithms, the SAC algorithm converges more rapidly and stably across all the three considered metrics. This effectiveness is primarily due to SAC’s ability to balance exploration and exploitation during policy updates, enhancing convergence efficiency. Consequently, SAC proves to be the most effective algorithm for achieving the goal of minimizing energy consumption, making it the best performer overall.
@@
In Fig. 7, we compare the trends in the loss function, the Top1 and Top5 classification accuracy of the global model of FLSimCo (without offloading) and our method (with offloading) under varying numbers of vehicles and local iterations per round. It is obvious that the loss function gradually decreases and converges, while the accuracy steadily increases and converges. Specifically, as illustrated in Fig. 7(a), when comparing the cases of 5 vehicles (with 5 models aggregated per round) and 10 vehicles (with 10 models aggregated per round) without task offloading, we observe that with 10 vehicles, the loss function fluctuates more and converges more slowly. This indicates that as the number of vehicles increases, the model diversity also increases, negatively impacting the effectiveness of federated aggregation. However, by offloading tasks to the RSU for training, we obtain models trained by the RSU, which are then sent back to the BS for aggregation along with the locally trained models. Therefore, with 5 vehicles and task offloading, the final number of models is doubled, resulting in 10 local models aggregated per round, which is the same as the number of models aggregated by the BS with 10 vehicles without task offloading. From the figure, we can see that the 5 vehicles with task offloading (10 models in total) not only outperform the 10 vehicles without task offloading but also the 5 vehicles without task offloading. This is because offloading some tasks to the RSU allows the completion of all scheduled iterations, and more local iterations help mitigate the negative impact of model diversity on federated aggregation. Fig. 7(b) and Fig. 7(c) show that in the first 60 episodes, the fewer models aggregated per round, the faster the accuracy improves. However, as the number of episodes increases, this diversity gradually diminishes, resulting in the accuracy of global with 5 and 10 vehicles without task offloading gradually converging. Additionally, when 5 vehicles participate in federated SSL and offload partial iterations task to RSU, the number of models used in the final aggregation is 10. During the first 60 episodes, the accuracy is not affected by model diversity. Instead, it increases rapidly and steadily. In the later stages of training, due to more iterations per round, the accuracy improvement is significantly better than others. This indicates that a greater number of iterations mitigates the negative effects of model diversity during aggregation.
@@
In Fig. 8, we compare the energy consumption and reward obtained by the SAC algorithm with different numbers of vehicles. From Fig. 8(a) and Fig. 8(b), it can be observed that the energy consumption curve shows a decreasing trend and gradually stabilize and the reward function shows a increasing trend and gradually stabilize. As the number of vehicles increases, the episodes required to reach a stable state also grow, reflecting the heightened complexity of task offloading and resource allocation in multi-vehicle scenarios. Overall, with more vehicles participating in training, total system energy consumption rises, while reward values decline. Additionally, these two figures clearly illustrate the SAC algorithm’s superiority in convergence speed and stability.